### v2

# EASSE-DE: Easier Automatic Sentence Simplification Evaluation for German

Abstract
In this work, we propose EASSE-multi, a framework for easier automatic sentence evaluation for languages other than English. Compared to the original EASSE framework, EASSE-multi does not focus only on English. It contains tokenizers and versions of text simplification evaluation metrics which are suitable for multiple languages. In this paper, we exemplify the usage of EASSE-multi for German TS resulting in EASSE-DE. Further, we compare text simplification results when evaluating with different language or tokenization settings of the metrics. Based on this, we formulate recommendations on how to make the evaluation of (German) TS models more transparent and better comparable. The code of EASSE-multi and its German specialisation (EASSE-DE) can be found in https://github.com/rstodden/easse-de.
[…]
EASSE-DE: Easier Automatic Sentence Simplification Evaluation
for German
Regina Stodden Department of Computational Linguistics Faculty of Arts and Humanities Heinrich Heine University Düsseldorf, Germany regina.stodden@hhu.de
1 Introduction
Automatic text simplification (TS) is a natural language processing (NLP) task that involves the development of algorithms and models to automatically transform complex textual content into more straightforward and accessible language. Manual or automatic evaluation is required to measure the quality of the generated simplifications. A good simplification should be grammatically correct, more simple and better readable than the original text and preserve the original meaning of it. For manual evaluation, people are asked to rate the extent of these three aspects for the generated simplification with respect to the original sentence. Because manual evaluation is very time-consuming, automatic metrics are used for a first quality check of sentence simplification models Alva-Manchego et al. (2020). Compared to manual evaluation methods, automatic evaluation methods facilitate a quick assessment of the output of various text simplification models, making it feasible to compare and iterate on different approaches efficiently. Further, with the increasing mass of evaluation data of different model approaches, it becomes challenging to evaluate this large number of generated texts manually. Automatic evaluation methods allow researchers to scale up their assessments to handle large datasets effectively Alva-Manchego et al. (2020).
Alva-Manchego et al. (2019) proposed an evaluation framework for easier automatic sentence simplification evaluation, called EASSE, to facilitate a comparison of TS models on existing test sets and on the same evaluation metrics as well as to unify the implementation of the evaluation metrics. EASSE is nowadays the common standard for evaluating English TS models. Although it is specified for only English TS evaluation, it is often also used to evaluate TS models of other languages, e.g., German (see, e.g., Trienes et al. 2022), Spanish (see, e.g., Gonzalez-Dios et al. 2022), French (see, e.g., Cardon and Grabar 2020), Spanish (see, e.g.,  Holmer and Rennes 2023) or on a multi-lingual benchmark Ryan et al. (2023). However, using EASSE on non-English texts raises some problems, e.g., the tokenizer is not adapted to the language of interest, the BERT-Score is evaluated on an English-only BERT model, and the readability scores are only designed for English.
In this paper, we present EASSE-multi, an adaptation of EASSE for languages other than English, to make the evaluation of non-English TS easier and more robust. We exemplify its usage for one language with several TS resources, i.e., German and the German EASSE variant, EASSE-DE. We further analyze the effects of different settings in EASSE-DE on TS metrics when evaluating non-English texts (i.e., German texts).
2 Related Work
2.1 Automatic Evaluation
In order to automatically evaluate text simplification, SARI Xu et al. (2016) is the primary metric to measure the overall simplicity quality. In more detail, SARI compares a generated simplification sentence with the source sentence and several references to estimate the quality of the lexical simplification. Further, most often BLEU Papineni et al. (2002) and BERT-Score-Precision Zhang* et al. (2020) are utilized to measure the similarity or meaning preservation between the original text and the system-generated simplification. Following Alva-Manchego et al. (2021), BERT-Score-Precision can also measure overall simplicity even if not implemented for this use case. Recently, the LENS score Maddela et al. (2023) has been proposed to measure the overall simplification quality of English simplifications; it is a trainable score trained on human assessments and English complex-simple pairs.
Readability formulas such as FRE or FKGL Flesch (1948) are also often used to estimate the readability of the system output Alva-Manchego et al. (2021). For a syntactical simplification evaluation, SAMSA Sulem et al. (2018b) has been proposed: SAMSA is a reference-less metric based on the annotation of semantic structures.
The reliability of these metrics for English TS evaluation has been questioned in research, e.g., see Sulem et al. (2018a), Tanprasert and Kauchak (2021), or Alva-Manchego et al. (2021). Another issue with automatic metrics is that the reliability of the scores has only been evaluated against human annotations of English annotations and that the correlations are not yet reproduced or repeated in other languages. To facilitate the assessment of the reliability of the metrics for German texts, we gather available system outputs and automatic scores regarding them in EASSE. As long as SARI, BLEU, and BERT-Score are still common practices in TS research, we will use them in our analysis but are also open to replacing or extending the metrics.
Furthermore, Stodden and Kallmeyer (2020) have shown that the way how English sentences are simplified differs from the German or Spanish ways. Hence, different simplification metrics might be required per language. An approach in this direction could be learnable metrics (per language) as LENS Maddela et al. (2023), BETS Zhao et al. (2023) or MeaningBERT Beauchemin et al. (2023), which are currently only applied to English texts.
2.2 Original EASSE Package
The original EASSE package Alva-Manchego et al. (2019) is designed to ease the automatic evaluation of English sentence simplification. It contains the implementation of automatic evaluation metrics, including SARI, BLEU, SAMSA, FKGL, and BERT-Score, as well as a linguistic feature analysis on the simplification pairs utilizing the TS-eval package by Martin et al. (2018). EASSE also stores English TS test sets and outputs of English TS systems, as well as builds an evaluation report regarding all specified metrics of all specified TS models to facilitate the whole evaluation process. It is commonly used to evaluate TS system outputs in English and other languages.
3 System Overview: EASSE-multi
In order to make EASSE language-independent and more robust for evaluating texts of languages other than English, we are proposing EASSE-multi (and its German variant EASSE-DE in the next section).
Therefore, we add a language constant to EASSE-multi to specify the currently evaluated language (e.g., “DE” for German in EASSE-DE). We also add SpaCy to the list of possible tokenizers to allow tokenization specified for languages other than English (see subsection 3.1).
The language constant also allows to choose language-specific evaluation metrics, e.g., readability metrics (see subsection 3.3), different models for BERT-Score (see subsection 3.2) and multi-lingual linguistic feature extraction (see subsection 3.4).
3.1 Tokenization
The original EASSE version currently supports 13a tokenization or white-space split tokenization (presuming pre-tokenized data). To include the language component into tokenization, we added the tokenizers of SpaCy Montani et al. (2023) and the extension Spacy-Stanza Qi et al. (2020) [^1] https://github.com/explosion/spacy-stanza as they currently support the tokenization of roughly 70 languages and also support linguistic annotations, e.g., part-of-speech tagging and dependency parsing, which will be relevant for the linguistic feature extraction.
3.2 Metrics
Evaluation metrics for TS are mostly language-independent, e.g., SARI, or BLEU, as they are n-gram-based methods. However, the n-grams depend on tokenization, which differs from language to language (see previous section). Following Zhang* et al. (2020), BERT-Score can be used for a specific language (e.g., using the English-only model RoBERTa (Liu et al., 2019)) or in a multi-lingual setting (e.g., using a multi-lingual model such as BERT-multilingual (Devlin et al., 2019)).
In EASSE-multi, the usage of the metrics is optimized regarding the evaluated language, as based on the language constant, the tokenizer and the BERT-model are chosen to fit non-English languages better.
3.3 Readability
Readability scores and the LENS-Score Maddela et al. (2023) are language-dependent, for the first due to included language-specific averages of word and sentence lengths and for the second due to training an evaluation score exclusively on English.
As an extension of EASSE, we also added readability formulas for languages other than English to EASSE-multi, which have already been implemented in the textstat package. For example, common readability scores for German are the Amstad’s adaption on the Flesch Reading Ease (FRE) or the Vienna non-fictional text formulas Bamberger and Vanecek (1984). LENS has not been reproduced for other languages; hence, it makes no sense to include it in EASSE-multi.
Following the criticism of Tanprasert and Kauchak (2021) regarding readability metrics for TS evaluation, we follow their recommendation and include average sentence length, number of syllables and number of splits in our report. Hence, we add these features to the default report.
3.4 Multi-lingual Feature Extraction
As argued in Tanprasert and Kauchak (2021) and Alva-Manchego et al. (2019), we include a few linguistic features to get more insights into the system-generated simplification. For this, we are using the feature extraction toolkit of the reference-less quality estimation tool (further called TS-eval) by Martin et al. (2018) for the English analysis and its extended language-independent version TS-eval-multi by Stodden and Kallmeyer (2020). We decided to use TS-eval-multi for feature extraction and not the similar language-independent feature extraction toolkit called LFTK Lee and Lee (2023) as both versions of TS-eval focus more on features for text simplification, whereas LFTK focuses more on features for readability assessment. The TS-eval package has also already been integrated into the evaluation package EASSE, which facilitates its extension to the multilingual TS-eval. Further, most of LFTK’s implemented features only apply to English. In future work, TS-eval-multi could be extended with features of LFTK. TS-eval-multi contains, for example, the parse tree height, cosine similarity between source and output based on pre-trained word embeddings, and length of phrases and clauses.
3.5 Additional Resources
The original EASSE framework also includes resources of English TS, i.e., English TS test sets, word lists, and system outputs of English TS models. With EASSE-multi, this component can be extended to the language of interest. We exemplify this with EASSE-DE and add only German resources (see section 4). However, the German resources can be easily replaced with resources of other languages.
3.6 Recommended Setting
At the moment, we cannot provide recommended settings per language except specifying the language constant, using SpaCy for tokenization, and using the multi-lingual BERT-Score. Further recommendations, for example, if case sensitivity is useful for the language of interest or determining which BERT version is more suitable for the language of interest, require more analysis which is out of the scope of this work. However, we recommend always naming which kind of settings have been used during evaluation as it can greatly influence the TS metrics. The settings should be reported in detail to ensure that the effect on the metric is due to the TS system and not the evaluation metrics’ settings.
Furthermore, it could be helpful to report the results of the baselines, e.g., src2src (source-to-source or complex-to-complex) or tgt2tgt (target-to-target or simple-to-simple). If the system outputs cannot be made available, it could help to verify on the gold data whether the applied evaluation method (e.g., in a replication experiment) is the same as the evaluation method used for an original experiment, as the results should be identical. Additionally, it could be helpful to re-evaluate the data comparing to. Therefore, we recommend making the system outputs publicly available (if the data is not restricted by license or copyright), e.g., as part of the EASSE-DE resources.
4 EASSE-DE: Using EASSE-multi for German TS Evaluation
We will exemplify the usage of EASSE-multi for one language, i.e., German, resulting in EASSE-DE. Therefore, we add German resources to EASSE-DE (see subsection 4.1), i.e., German sentence simplification test sets (see subsubsection 4.1.1)), and available outputs of German TS systems regarding these test sets (see subsubsection 4.1.2). Further, we analyze whether and to what extent differences exist when evaluating German text simplification with the original evaluation framework EASSE or its adaptation EASSE-DE (see subsection 4.2).
4.1 German TS Resources
4.1.1 German TS Test Sets
For a better overview of available test sets for German sentence simplification, we have added gold data, i.e., manually simplified complex-simple sentence pairs, to EASSE-DE. In more detail, EASSE-DE refers to nine test sets, i.e., ABGBMeister (2023), APA-LHA-OR-A2 Spring et al. (2021), APA-LHA-OR-B1 Spring et al. (2021), BiSECT Kim et al. (2021), DEplain-APA Stodden et al. (2023), DEplain-web Stodden et al. (2023), TextComplexityDE Naderi et al. (2019), GEOlino Mallinson et al. (2020), and Simple-German-Corpus Toborek et al. (2023). We refer to Table 1 for more meta data of the test sets.
complex simple
name target group domain size # ref. n:m FREnormal-$\downarrow$ sent. len.normal-$\uparrow$ word len.normal-$\uparrow$ FREnormal-$\uparrow$ sent. len.normal-$\downarrow$ word len.normal-$\downarrow$
ABGB non-experts law 448 2 40% 42.75 24.85 1.83 44.6 22.39 1.89
APA_LHA-or-a2 Non-native speaker news 500 1 6 % 44.7 20.2 1.92 69.55 11.27 1.78
APA_LHA-or-b1 Non-native speaker news 500 1 8 % 43.7 20.48 1.93 62.6 12.82 1.83
BiSECT people w. reading problems politics 753 1 100 % 8.55 30.24 2.01 35.85 15.72 1.98
DEplain-APA Non-native speaker news 1,231 1 27 % 58.75 11.92 1.86 65.8 10.55 1.79
DEplain-web mixed web/mixed 1,846 1 57 % 62.95 19.13 1.64 77.9 10.76 1.57
GEOlino children encyclopedia 663 1 40 % 61.5 13.31 1.7 66.0 9.94 1.66
simple-german-corpus mixed web/mixed 391 1 73 % 41.15 13.96 2.0 65.4 9.31 1.83
TextComplexityDE Non-native speaker encyclopedia 250 1 83 % 28.1 27.75 2.08 51.2 14.17 1.9
Table 1: Overview Test Sets for German Sentence Simplification which are included in EASSE-DE. Including the target group, domain, size in sentence pairs, number of references, percentage of $n:m$ alignments, word length measured in syllables, and sentence length measured in words.
4.1.2 German TS Models
For German TS, a few models are available or reproducible, e.g., ZEST, by Mallinson et al. (2020), sockeye by Spring et al. (2021), custom-decoder-ats by Anschütz et al. (2023), the few-shot approaches on BLOOM by Ryan et al. (2023), or the mBART models by Stodden et al. (2023). A more detailed description and analysis of German TS models, including their reproduction, has been recently proposed by Stodden (2024 (to appear). The system outputs of all reproduced German TS models (see Table 2) have been added to EASSE-DE Table 2 to facilitate a better comparison between existing models and models which will be newly proposed in future.
System Name Reference Type Training Data # Simp. Pairs URL
hda-etr Siegel et al. (2019) rule-based - -
https://github.com/hdaSprachtechnologie/easy-to-understand_language
sockeye-APA-LHA
Spring et al. (2021) &
Ebling et al. (2022)
seq2seq
APA-LHA OR-A2 &
APA-LHA OR-B1
8,455 &
9,268
https://github.com/ZurichNLP/RANLP2021-German-ATS
sockeye-DEplain-APA - seq2seq DEplain-APA 10,660
https://huggingface.co/DEplain
mBART-DEplain-APA Stodden et al. (2023)
fine-tuned
seq2seq
DEplain-APA 10,660
https://huggingface.co/DEplain/trimmed_mbart_sents_apa
mBART-DEplain-APA+web Stodden et al. (2023)
fine-tuned
seq2seq
DEplain-APA+web
10,660 +
1,594
https://huggingface.co/DEplain/trimmed_mbart_sents_apa_web
mT5-DEplain-APA -
fine-tuned
seq2seq
DEplain-APA 10,660
https://huggingface.co/DEplain
mT5-SGC -
fine-tuned
seq2seq
SGC 4,430
https://huggingface.co/DEplain
BLOOM-zero Ryan et al. (2023)
zero-shot
AR model
https://github.com/XenonMolecule/MultiSim
BLOOM-sim-10 Ryan et al. (2023)
few-shot
AR model
TCDE19 &
GEOlino
200 & 959
https://github.com/XenonMolecule/MultiSim
BLOOM-random 10 Ryan et al. (2023)
few-shot
AR model
TCDE19 &
GEOlino
200 & 959
https://github.com/XenonMolecule/MultiSim
custom-decoder-ats Anschütz et al. (2023)
AR model +
fine-tuned
seq2seq
Simplified, monolingual
German data &
20Minuten
544,467 &
17,905
https://huggingface.co/josh-oo/custom-decoder-ats
Table 2: Overview of German TS models including training details (i.e., training data and size of training samples). Each line separates different model types. Adaptation from Stodden (2024 (to appear).
4.2 Comparison of EASSE and EASSE-DE
In the following section, we present and analyse the metric scores when using EASSE or EASSE-DE, including different settings on three test sets of one TS model.
4.2.1 Method
Evaluation Settings.
In the comparative analysis, we focus on the settings in EASSE regarding i) language specification (i.e., English vs. German), ii) tokenization method (i.e., none vs 13a vs SpaCy), iii) BERT model version (i.e., RoBERTa-large vs BERT-base-multilingual-cased), iv) FRE version (English vs German). Due to their n-gram-based approach, we expect the tokenization method to have an effect on SARI and BLEU but not on BERT-Score-Precision.
German TS Test Sets.
In the analysis, we evaluate on three available German TS test sets: DEplain-APA (Stodden et al., 2023), DEplain-web (Stodden et al., 2023), and TextComplexityDE (Naderi et al., 2019). These test sets are all manually simplified and manually aligned, and, therefore, we expect a higher simplification quality for them as for other test sets, e.g., BiSECT Kim et al. (2021) [^2] BiSECT is generated using machine translation of English texts. Due to this augmentation strategy, the German version includes encoding errors. or APA-LHA Spring et al. (2021) [^3] The training and validation sets of APA-LHA are automatically aligned, and, hence, more faulty compared to manually aligned corpora.. Further, these three test sets include texts of different domains (news, web, and Wikipedia), and their simplification addresses different target groups (non-native speakers and people with cognitive disabilities). Hence, they represent different kinds of simplifications and therefore seem to be a good choice for our analysis.
German TS Model.
Further, we have selected the generated simplifications of one model, i.e., mBART-DEplain-APA+web. Reasons for the choice of this model are that it is ready-to-use without additional examples, and, following Stodden (2024 (to appear), this model achieves the best BERT-Scores across several test sets. In comparison, the BLOOM models by Ryan et al. (2023) are few-shot models that require additional complex-simple pairs to generate simplifications.
4.2.2 Results
The results of the mBART-APA+web model with different settings are presented in Table 3.444To ensure that the effects are not due to the system but to the evaluation changes, we also add the results of the identity baseline (see Table 4). In the following, we analyse the differences regarding tokenization, readability scores, multi-lingual BERT-Score, and multi-lingual feature extraction.
Tok. Lang. BLEUnormal-$\uparrow$ SARInormal-$\uparrow$ BS-Pnormal-$\uparrow$ FREnormal-$\uparrow$
TCDE19 (n = 250) spacy EN 18.56 37.69 0.39 57.37
spacy DE 17.75 37.37 0.55 43.65
13a DE 18.04 37.41 0.55 43.55
none DE 16.04 37.47 0.55 43.65
DEplain-APA (n = 1231) spacy EN 30.59 34.79 0.48 78.25
spacy DE 28.03 33.81 0.64 65.2
13a DE 28.37 33.92 0.64 65.2
none DE 24.69 32.88 0.64 65.2
DEplain-web (n = 1846) spacy EN 18.37 34.21 0.27 76.52
spacy DE 17.99 34.07 0.44 69.05
13a DE 18.17 34.10 0.44 69.05
none DE 15.97 33.67 0.44 69.05
Table 3: Scores of trimmed-mbart-DEplain-APA+web when using different language settings and tokenizers.
Tokenization.
As expected, different tokenization methods (including language specification) affect the calculation of metrics used for TS evaluation. The last three rows in each block of Table 3 show the differences in the scores when using different tokenization strategies. We can see that the BERT-score is always the same for all settings due to the sub-word tokenization in BERT.
The FRE scores are also robust across all test sets when looking at the trimmed-mBART results, but in Appendix A Table 4, we see slightly more differences. The SARI scores also change slightly, i.e., to less than 1 point in all settings, whereas the differences in the BLEU scores range between 2 to 3 points in all test sets. In conclusion, when comparing one model against another with a slightly different evaluation setting (here, the tokenizer), even these small changes can be wrongly interpreted as an improvement of the model idiosyncrasy. However, it is only due to the different settings. Therefore, we recommend stating all settings chosen for evaluation for a more reliable comparison between models.
Readability Metrics.
As can be seen in Table 3, the scores are quite different wrt. to FRE for the English and German settings (see first two rows in each block). The results are different due to the different constants of the formulas and their dependency on different tokenization and syllable splitting. When interpreting the readability scores, they also result in different categories: Following Amstad (1978), the simplifications with the English setting on DEplain-APA and DEplain-web can be described as “ease” whereas they are categorized as “simple” using the German setting. In summary, the language adaptation of readability scores can make a noticeable difference when interpreting the simplification results.
BERT-Score.
As shown in the first two rows of each row-block in Table 3, changing the transformer model of the BERT-Score significantly affects the BERT-Score. The scores using the multi-lingual model are much higher than those using the only-English model. Hence, the choice of the BERT model seems to have a high effect on the TS evaluation.
Linguistic Feature Extraction.
Due to space restrictions, we do not include analysis regarding linguistic feature extraction. We refer the interested reader to the system reports in our code repository [^5] https://github.com/rstodden/easse-de, which also include the feature extraction results.
5 Discussion & Conclusion
We have proposed EASSE-multi, which facilitates easy evaluation of sentence simplification in multiple languages. Therefore, we have extended the original EASSE package with a language-constant tokenizer, language-dependent version of BERT-Score, and language-wise readability scores.
Further, we have exemplified using EASSE-multi for German TS evaluation in the form of EASSE-DE. In comparing the results generated by EASSE and EASSE-DE, we have shown that it is important to consider the text’s language when evaluating. Following that, we recommend using EASSE-DE over EASSE when evaluating German sentence simplification models as it includes language-sensitive evaluation metrics. Even if the scores per metric might be lower when using EASSE-DE than EASSE, we argue that these are more reliable due to the language-sensitive metrics.
Further, we argue that it is unreliable to compare scores (maybe originating from different papers) as they might be generated by using different evaluation settings. Before making a comparison, we recommend verifying whether the same settings of the metric have been used in both experiments (the referenced and the new one). Otherwise, the differences in the scores might not be dependent on the model changes (which is the question of interest) but on, for example, different kinds of tokenization. Therefore, we strongly recommend always specifying the settings or, even better, the implementation of the metrics used for the evaluation, as it can have a huge impact on the reported scores. We identified the following aspects which should be reported accompanied with automatic evaluation: 1. language setting (e.g., EN, or DE) for features (e.g., BERT-Score, FRE, or word length), 2. tokenizer (e.g., none, 13a, or SpaCy), 3. lower casing (True or False), 4. BERT-Score model (e.g., RoBERTA-large, mT5, or BERT-base-multilingual-cased)
6 Future Work
Even if most of the scores are language-independent or can be easily adapted to work for other languages, as shown previously, there still might be problems in using the same scores for different languages due to language idiosyncrasies and different simplification operations per language. Approaches in the direction of language-wise evaluation of non-English TS could be learnable metrics (per language) as already proposed for English, e.g., LENS, BETS, or MeaningBERT. In future work, we want to investigate learnable metrics for non-English languages to fit the language idiosyncrasies better and add them to EASSE-DE.
Further, we would like to extend EASSE-DE to include more German TS resources. We hope that EASSE-DE will be useful for German TS researchers and invite them to contribute their test sets or system outputs to EASSE-DE.
Appendix A Results of Identity Baseline
Tok. Lang. BLEUnormal-$\uparrow$ SARInormal-$\uparrow$ BS-Pnormal-$\uparrow$ FREnormal-$\uparrow$
TCDE19 (n = 250) spacy EN 28.22 15.31 0.37 39.16
spacy DE 27.31 14.99 0.55 28.1
13a DE 27.49 15.05 0.55 28.0
none DE 24.43 13.78 0.55 28.1
DEplain-APA (n = 1231) spacy EN 29.28 16.17 0.45 77.64
spacy DE 26.89 15.25 0.63 58.75
13a DE 7.25 15.35 0.63 64.6
none DE 23.33 13.75 0.63 58.75
DEplain-web (n = 1846) spacy EN 21.24 12.09 0.25 70.33
spacy DE 20.85 11.93 0.42 62.95
13a DE 20.89 11.94 0.42 62.95
none DE 18.82 10.9 0.42 62.95
Table 4: Scores of identity baseline on three test sets when using different language settings and tokenizers.